### v3

# Barrow Cosmology and Big-Bang Nucleosynthesis

Abstract
Using thermodynamics-gravity conjecture, we present the formal derivation of the modified Friedmann equations inspired by the Barrow entropy, $S\sim A^{1+\delta/2}$, where $0\leq\delta\leq 1$ is the Barrow exponent and $A$ is the horizon area. We then constrain the exponent $\delta$ by using Big-Bang Nucleosynthesis (BBN) observational data. In order to impose the upper bound on the Barrow exponent $\delta$, we set the observational bound on $\left|\frac{\delta T_{f}}{T_{f}}\right|$. We find out that the Barrow parameter $\delta$ should be around $\delta\simeq 0.01$ in order not to spoil the BBN era. Next we derive the bound on the Barrow exponent $\delta$ in a different approach in which we analyze the effects of Barrow cosmology on the primordial abundances of light elements i.e. Helium ${}^{4}\textit{He}$, Deuterium $D$ and Lithium ${}^{7}\textit{Li}$. We observe that the deviation from standard Bekenstein-Hawking expression is small as expected. Additionally we present the relation between cosmic time $t$ and temperature $T$ in the context of modified Barrow cosmology. We confirm that the temperature of the early universe increases as the Barrow exponent $\delta$ (fractal structure of the horizon) increases, too.
I Introduction
After the discovery of the black hole thermodynamics as well as four laws of black hole mechanics Bardeen ; Bekenstein ; Hawking , the profound connection between gravity and thermodynamics has been speculated. In particular, the derivation of the Einstein field equation of gravity from the first law of thermodynamics by Jacobson Jacobson , reveals that gravity is nothing but the thermodynamics of spacetime at large scales. The deep connection between gravity and thermodynamics has been confirmed in different setups Verlinde ; Padman ; Padmann . In the background of Friedmann-Robertson-Walker (FRW) universe, it has been shown that Friedmann equations describing the evolution of the universe can be translated to the first law of thermodynamics on the apparent horizon Eling ; Akbar ; MAkbar ; Cai ; Asheykhi ; Sheykh . The studies were also generalized to the modified cosmological models, based on the modification of the Bekenstein-Hawking area law of entropy associated with the horizon Paranj ; Jamil ; Caiii ; Wang ; Setare ; Fan ; Tsheykhi .
An interesting modification to the area law was proposed by J. D. Barrow, who argued that black hole horizon may have a fractal structure due to quantum fluctuations JohnD . In this viewpoint, quantum-gravitational effects may deform the black hole horizon which can be described by a fractal structure as a first approximation JohnD . Consequently the black hole entropy gets a deviation from the standard Bekenstein-Hawking area law, and is modified as John
$S_{h}=\left(\frac{A}{A_{0}}\right)^{1+\delta/2},$ (1)
where $A$ is the horizon area and $A_{0}$ is the Planck area. The exponent $\delta$ ranges in the interval $0\leq\delta\leq 1$, and it shows the amount of quantum-gravitational deformation effects John ; HawkingS ; Hooft . The extreme values $\delta=1$ stands for the most deformed and intricate structure, while for $\delta=0$, the area law of entropy will be restored ($A_{0}\rightarrow 4G$) Shey1 . In the framework of thermodynamics-gravity conjecture, it was argued that one can derive the friedmann equations of FRW universe by applying the first law of thermodynamics on the apparent horizon. In this derivation, the entropy expression plays a crucial role. Any modification to the area law, shall modify the corresponding Friedmann equations. Therefore, by applying the first law of thermodynamics on the apparent horizon, and using the modified Barrow entropy (1) instead of Bekenstin-Hawking one, one can get the modified Friedmann equations Shey1 which leads to a so-called modified Barrow cosmology Shey2 . The Barrow exponent $\delta$ con be constrained by using observational data. For example, using BBN data one can impose constraints on the exponent of Barrow entropy John . It was shown that the Barrow exponent should be $\delta<1.4\times 10^{-4}$ in order not to spoil the BBN epoch. This implies that the deviation from standard area law of entropy should be small as expected John .
BBN refers to the process that took place in the early universe, shortly after the Big Bang, when the universe was extremely hot and dense. BBN occurred roughly between the very first centiseconds after the Big-Bang ($\sim 0.01$ sec) and a few hundred seconds ($\sim 3$ minutes) after it. During BBN, the temperature of the universe allows for nuclear reactions to occur, leading to the formation of light elements such as $H$, $He$ and $Li$. As the universe expanded and cooled, these reactions slowed down and eventually stopped, resulting in the abundance of these light elements observed today. Hence BBN, together with cosmic microwave background radiation provide a strong evidence of the hot, dense state of the early universe. In general, the principles of BBN allows to impose strict constraints on various cosmological models. The main idea is to demand consistency between theoretical predictions, and BBN observational data. In this regard, the deviation of the freeze-out temperature in comparison to standard cosmological model can be calculated. This allows to compare the theoretical results of the modified cosmological models with the observational constraints on $\left|\frac{\delta T_{f}}{T_{f}}\right|$John ; Luciano ; Anish . A modified cosmological model can also affect the primordial abundances of aforementioned light elements in the early universe Anish . It is important to note that although the standard BBN theory successfully explains the observed amounts of elements like Helium and Hydrogen, it encounters significant issues when it comes to Lithium Boran . This leads to what is known as the cosmological Lithium problem, which is a highly debated topic in the modern cosmology. It is interesting to check whether this problem can be addressed in the framework of the modified cosmology or not. The aim is to investigate the the cosmological Lithium problem in the background of the modified cosmology, to check under which conditions and which values for the model parameter, it can be alleviated.
Our aim in this paper is to explore the implications of Barrow cosmology on the BBN and derive constraints on Barrow exponent $\delta$. In this regards, we can check the viability of the Barrow cosmology by imposing the BBN conditions. The most important difference between this work and the approach presented in John is that we modify the left-hand side of the Friedmann equations (the geometry side), while in John the authors modify Friedmann equations in such a way that the Barrow entropy correction appears as an additional energy density $\rho_{de}$ and pressure $p_{de}$ which resemble dark energy components in the right-hand side of the Friedmann equations. Additionally, in deriving the modified Friedmann equations from first law of thermodynamics, we neglect the integration constant, while in John the integration constant is interpreted as the cosmological constant which influences the dynamics of the universe. Furthermore, we should also note that different approximations were employed in our approach compared to that in John . We also present the relation between cosmic time $t$ and temperature $T$ in the early universe and in the context of the modified Barrow cosmology which is novel and has not been addressed in the literature. We will see that the temperature of the early universe increases as the fractal structure of the boundary of our universe increases.
The structure of the paper is as follows. In the next section, we apply the first law of thermodynamics, $dE=T_{h}dS_{h}+WdV$ on the apparent horizon, and construct the modified Friedmann equations based on Barrow entropy. In section III, we use BBN data in the background of the modified barrow cosmology to constrain the Barrow exponent $\delta$. We also present the relation between time and temperature in the context of Barrow cosmology in this section. We finish with closing remarks in section V.
II Barrow entropy and Modified Friedman Equations
Let us start with a homogeneous and isotropic FRW universe with line elements
$ds^{2}={h}_{\mu\nu}dx^{\mu}dx^{\nu}+\tilde{r}^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (2)
where $\tilde{r}=a(t)r$, $x^{0}=t,x^{1}=r$, and $h_{\mu\nu}$=diag $(-1,a^{2}/(1-kr^{2}))$ represents the two dimensional metric. The open, flat, and closed universes corresponds to $k=0,1,-1$, respectively. The radius of the apparent horizon, which is a thermodynamically suitable horizon for the universe, is given by Sheyem
$\tilde{r}_{A}=\frac{1}{\sqrt{H^{2}+k/a^{2}}}.$ (3)
The associated temperature with the apparent horizon is defined as MAkbar ; Sheyem
$T_{h}=-\frac{1}{2\pi\tilde{r}_{A}}\left(1-\frac{\dot{\tilde{r}}_{A}}{2H\tilde{r}_{A}}\right).$ (4)
For $\dot{\tilde{r}}_{A}\leq 2H\tilde{r}_{A}$, the temperature becomes negative which is not physically acceptable and hence one should define $T=|\kappa|/2\pi$. Note that within an infinitesimal internal of time $dt$ one may assume $\dot{\tilde{r}}_{A}\ll 2H\tilde{r}_{A}$, which physically means that the apparent horizon radius is fixed. Thus there is no volume change in it and one may define $T=1/(2\pi\tilde{r}_{A})$ CaiKim . The relation between horizon temperature and Hawking radiation has been confirmed cao .
We further assume the energy content of the Universe is in the form of a perfect fluid with energy-momentum tensor,
$T_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu},$ (5)
where $\rho$ and $p$ are the energy density and pressure, respectively and we assume the conservation of energy holds, namely, $\nabla_{\mu}T^{\mu\nu}=0$. In the background of FRW universe, this leads to
$\dot{\rho}+3H(\rho+p)=0,$ (6)
where $H=\dot{a}/a$ is the Hubble parameter. The work density associated with the volume change, in an expanding universe, is given by Hay2
$W=-\frac{1}{2}T^{\mu\nu}h_{\mu\nu},$ (7)
where in the background of FRW background can be calculated as
$W=\frac{1}{2}(\rho-p).$ (8)
We also assume the first law of thermodynamics on the apparent horizon is satisfied and has the form
$dE=T_{h}dS_{h}+WdV,$ (9)
where $E=\rho V$ is the total energy of the Universe enclosed by the apparent horizon, and $T_{h}$ and $S_{h}$ are, respectively, the temperature and entropy associated with the apparent horizon. Differentiating $E=\rho V$, yields
$dE=4\pi\tilde{r}_{A}^{2}\rho d\tilde{r}_{A}+\frac{4\pi}{3}\tilde{r}_{A}^{3}\dot{\rho}dt.$ (10)
where $V=\frac{4\pi}{3}\tilde{r}_{A}^{3}$ is the volume enveloped by a 3-dimensional sphere with the area of apparent horizon $A=4\pi\tilde{r}_{A}^{2}$. Combining with the conservation equation (6), we get
$dE=4\pi\tilde{r}_{A}^{2}\rho d\tilde{r}_{A}-4\pi H\tilde{r}_{A}^{3}(\rho+p)dt.$ (11)
Taking differential form of Barrow entropy (1), we find
$dS_{h}$ $=$ $(2+\delta)\left(\frac{4\pi}{A_{0}}\right)^{1+\delta/2}{\tilde{r}_{A}}^{1+\delta}\dot{\tilde{r}}_{A}dt.$ (12)
Inserting relation (8), (11) and (12) in the first law of thermodynamics (9) and using definition (4) for the temperature, after some calculations, we find the differential form of the Friedmann equation as
$\frac{2+\delta}{2\pi A_{0}}\left(\frac{4\pi}{A_{0}}\right)^{\delta/2}\frac{d\tilde{r}_{A}}{\tilde{r}_{A}^{3-\delta}}=H(\rho+p)dt.$ (13)
Combining with the continuity equation (6), arrive at
$-\frac{2+\delta}{2\pi A_{0}}\left(\frac{4\pi}{A_{0}}\right)^{\delta/2}\frac{d\tilde{r}_{A}}{\tilde{r}_{A}^{3-\delta}}=\frac{1}{3}d\rho.$ (14)
Integration yields
$\frac{2+\delta}{2-\delta}\left(\frac{4\pi}{A_{0}}\right)^{\delta/2}\frac{1}{2\pi A_{0}}\frac{1}{\tilde{r}_{A}^{2-\delta}}=\frac{\rho}{3},$ (15)
where we have set the integration constant equal to zero. Substituting $\tilde{r}_{A}$ from Eq. (3) we immediately arrive at
$\frac{2+\delta}{2-\delta}\left(\frac{4\pi}{{A_{0}}}\right)^{\delta/2}\frac{1}{2\pi A_{0}}\left(H^{2}+\frac{k}{a^{2}}\right)^{1-\delta/2}=\frac{\rho}{3}.$ (16)
The above equation can be further rewritten as
$\left(H^{2}+\frac{k}{a^{2}}\right)^{1-\delta/2}=\frac{8\pi G_{\rm eff}}{3}\rho,$ (17)
where the effective Newtonian gravitational constant is defined as
$G_{\rm eff}\equiv\frac{A_{0}}{4}\left(\frac{2-\delta}{2+\delta}\right)\left(\frac{A_{0}}{4\pi}\right)^{\delta/2}.$ (18)
In this way we obtain the modified Friedmann equation inspired by Barrow entropy. Thus, starting from the first law of thermodynamics at the apparent horizon of a FRW universe, and assuming that the apparent horizon area has a fractal features, due to the quantum-gravitational effects, we derive the corresponding modified Friedmann equation of FRW universe with any spatial curvature. It is important to note that in this case, there is no dark energy component emerged from the Barrow entropy that was derived in Ref. Emm2 . In the limiting case where $\delta=0$, the area law of entropy is recovered and we have $A_{0}\rightarrow 4G$. In this case, $G_{\rm eff}\rightarrow G$, and Eq. (17) reduces to the standard Friedmann equation in Einstein gravity.
III Constraints on Barrow Exponent inspired by BBN
In this section we examine the BBN in the framework of Barrow cosmology to constrain the Barrow exponent $\delta$.
III.1 BBN in Barrow cosmology
The modified Friedmann equation inspired by the Barrow entropy in a flat universe ($k=0$), can be written as
$H=\left(\frac{8\pi G_{\rm eff}}{3}\rho\right)^{\frac{1}{2-\delta}}.$ (19)
Substituting $G_{\rm eff}$ from (18) in to above equation, we get
$H=\left(\frac{8\pi\rho}{3}\right)^{\frac{1}{2-\delta}}\left[\left(\frac{A_{0}}{4}\right)\left(\frac{2-\delta}{2+\delta}\right)\left(\frac{A_{0}}{4\pi}\right)^{(\delta/2)}\right]^{\frac{1}{2-\delta}}.$ (20)
We can rewrite (19) in the form
$H(T)\equiv Z(T)H_{GR}(T),$ (21)
where $H_{GR}=\sqrt{\frac{8\pi}{3M_{p}^{2}}\rho(T)}$ is the Hubble function in the standard cosmology and Z(T)is Amplication factor. In the limiting case where $\delta=0$ we have $A_{0}\rightarrow 4G$, and (20) can be rewritten in the form
$H(T)=\left(\frac{8\pi\rho}{3}\right)^{\frac{1}{2-\delta}}\left[G\;\left(\frac{2-\delta}{2+\delta}\right)\left(\frac{G}{\pi}\right)^{\delta/2}\right]^{\frac{1}{2-\delta}},$ (22)
and $Z(T)$ is defined as
$Z(T)\equiv\frac{H}{H_{GR}(T)}\rightarrow Z(T)\equiv$
$\left[\left(\frac{8\pi}{3M_{p}^{2}}\rho\right)\left(\frac{2-\delta}{2+\delta}\right)\left(\frac{1}{\pi M_{p}^{2}}\right)^{\delta/2}\right]^{\frac{1}{2-\delta}}\left(\frac{8\pi}{3M_{p}^{2}}\rho\right)^{-1/2}$
$=\left[\left(\frac{8\pi}{3M_{p}^{2}}\rho\right)^{\delta/2}\left(\frac{2-\delta}{2+\delta}\right)\left(\frac{1}{\pi M_{p}^{2}}\right)^{\delta/2}\right]^{\frac{1}{2-\delta}}.$ (23)
Inserting energy density of relativistic particles which is given by $\rho(T)=\frac{\pi^{2}}{30}g_{*}T^{4}$ ( $g_{*}\sim 10$ is the effective number of degrees of freedom and $T$ is the temperature) into Eq. (III.1) and using $(G)^{-1}=M_{p}^{2}$ (in the units where $\hbar=c=k_{B}=1$), we find the amplification factor $Z(T)$ as
$Z(T)\equiv\left(\frac{T}{M_{p}}\right)^{\frac{2\delta}{2-\delta}}\left[\frac{2\pi}{3}\sqrt{\frac{g_{*}}{5}}\left(\frac{2-\delta}{2+\delta}\right)^{1/{\delta}}\right]^{\frac{\delta}{2-\delta}}\equiv\eta\left(\frac{T}{M_{P}}\right)^{\nu},$ (24)
where we have defined
$\eta\equiv\left[\frac{2\pi}{3}\sqrt{\frac{g_{*}}{5}}\left(\frac{2-\delta}{2+\delta}\right)^{1/{\delta}}\right]^{\frac{\delta}{2-\delta}}\;,\;\;\nu\equiv\frac{2\delta}{2-\delta}.$ (25)
As expected, for $\delta=0$ (which implies that $\nu=0$), the amplification factor takes the value $Z(T)=1$, and Einstein gravity is recovered.
Now we examine the BBN through Barrow cosmology. The primordial ${}^{4}\textit{He}$ formation occurs at a time when the temperature was about $T\sim 100MeV$ and the constituents of the energy and number density at this time were electrons, positrons, neutrinos and antineutrinos, plus photons. All these particles are in thermal equilibrium because of their rapid collisions. Protons and neutrons were kept in thermal equilibrium through the interactions
$\nu_{e}+n\leftrightarrow p+e^{-},$
$e^{+}+n\leftrightarrow p+\overline{\nu}_{e},$ (26)
$n\leftrightarrow p+\overline{\nu}_{e}+e^{-}.$
The neutron abundance can be estimated through calculating the conversion rate of neutrons to protons, indicated with $\lambda_{np}(T)$ and its inverse $\lambda_{pn}(T)$. Hence the weak interaction rates at enough high temperature is
$\lambda_{tot}(T)=\lambda_{np}(T)+\lambda_{pn}(T),$ (27)
where $\lambda_{np}(T)$ is given by the sum of (III.1) rates, i.e.,
$\lambda_{np}(T)$ $=$ $\lambda_{(n+\nu_{e}\rightarrow p+e^{-})}+\lambda_{(n+e^{+}\rightarrow p+\bar{\nu}_{e})}$ (28)
$+\lambda_{(n\rightarrow p+e^{-}+\bar{\nu}_{e})}.$
While the relation between the rates $\lambda_{np}(T)$ and $\lambda_{pn}(T)$ is $\lambda_{np}(T)=e^{{\mathcal{Q}}/T}\lambda_{pn}(T)$, with $\mathcal{Q}=m_{n}-m_{p}$ the neutron and proton mass difference. From Eq. (27) one can obtain JohnD ; Bernstein ; Turner
$\lambda_{tot}(T)=4AT^{3}(4!T^{2}+2\times 3!\mathcal{Q}T+2!\mathcal{Q}^{2})$
$\simeq qT^{5}+\mathcal{O}\left(\frac{\mathcal{Q}}{T}\right),$ (29)
with $A=1.02\times 10^{-3}GeV$ and $q=9.6\times 10^{-10}GeV^{-4}$.
We can estimate the primordial mass fraction of ${}^{4}\textrm{He}$, by using Turner
$Y_{p}\equiv\lambda\frac{2x(t_{f})}{1+x(t_{f})},$ (30)
where $\lambda=e^{-(t_{n}-t_{f})/\tau}$, with $t_{f}$ the weak interactions freeze-out time, $t_{n}$ the nucleosynthesis freeze-out time, $\tau=\lambda^{-1}_{(n\rightarrow p+e^{-}+\overline{\nu}_{e})}\simeq 887sec$ the neutron mean lifetime Olive , and $x(t_{f})=e^{-\mathcal{Q}/T(t_{f})}$ is the equilibrium ratio of neutron to proton. The function $\lambda(t_{f})$ is the fraction of neutrons decaying into protons inside the range $t\in[t_{f},t_{n}]$. In any modified cosmology, the Hubble parameter $H$ will have a deviation from $H_{GR}$, and hence $T_{f}$ will also deviate from the GR results. Consequently, the primordial mass fraction $Y_{p}$ will also present a deviation $\delta Y_{p}$ from the standard cosmological model. From (30) we have
$lnY_{p}=$ $ln\lambda+ln2+lnX(t_{f})-ln(1+X(t_{f}))$
$\rightarrow\;\frac{\delta lnY_{p}}{\delta T_{f}}=\frac{\delta Y_{p}}{Y_{p}\delta T_{f}}=\frac{\delta}{\delta T_{f}}\left[-\frac{t_{n}-t_{f}}{\tau}\right.$
$\left.+ln2-\frac{\mathcal{Q}}{T(t_{f})}-ln(1+X(t_{f})\right]$
$=$ $-\frac{2}{\tau T_{f}^{3}}+\frac{\mathcal{Q}}{T_{f}^{2}}-\frac{\delta X(t_{f})}{\delta T_{f}(1+X(t_{f}))}$
$=-\frac{2t_{f}}{T_{f}\tau}+\frac{\mathcal{Q}}{T_{f}^{2}}-\frac{X(T_{f})}{X(T_{f})+1}\,\frac{\mathcal{Q}}{T_{f}^{2}}.$ (31)
Thus, we get
$\frac{\delta Y_{p}}{Y_{p}\delta T_{f}}=$ $-\frac{2t_{f}}{T_{f}\tau}+\frac{\mathcal{Q}}{T_{f}^{2}}\left(1-\frac{Y_{p}}{2\lambda}\right)$
$=\frac{1}{T_{f}}\left[-\frac{2t_{f}}{\tau}+\frac{\mathcal{Q}}{T_{f}}\left(1-\frac{Y_{p}}{2\lambda}\right)\right].$ (32)
And hence deviations from the primordial mass fraction $\delta Y_{p}$ can be written as
$\delta Y_{p}=Y_{p}\left[\left(1-\frac{Y_{p}}{2\lambda}\right)ln\left(\frac{2\lambda}{Y_{p}}-1\right)-\frac{2t_{f}}{\tau}\right]\frac{\delta T_{f}}{T_{f}}.$ (33)
Here relations $t_{f}\simeq T^{-2}$ (because BBN happens at the radiation dominated epoch), $ln\left(\frac{1}{X(t_{f})}\right)=ln\left(\frac{2\lambda}{Y_{p}}-1\right)$ and $\frac{X}{X+1}=\frac{Y_{p}}{2\lambda}$ have been used and we set $\delta T(t_{n})=0$ for the reason that $T_{n}$ is fixed by the deuterium binding energy Torres ; Lambiase ; Lambiasee ; Lambiaseee ; capoz .
The observational estimations of the fractional mass $Y_{p}$ of ${}^{4}\textit{He}$ are Aver
$Y_{p}=0.2449\pm 0.0040\;,\;\;\;|\delta Y_{p}|<10^{-4}.$ (34)
Inserting these into Eq. (33), the upper bound on ${\delta T_{f}}/{T_{f}}$ will be extracted as
$\left|\frac{\delta T_{f}}{T_{f}}\right|<4.7\times 10^{-4}.$ (35)
The freeze-out temperature $T_{f}$ is the temperature at which particles decouple, and corresponds to the time where $H=\lambda_{tot}(T)=qT^{5}$, with $H$ is given by (19), and hence we have
$H=$ $\sqrt{\frac{8\pi}{3M_{p}^{2}}\left(\frac{\pi^{2}g_{*}}{30}T_{f}^{4}\right)}\;\times\;\eta\left(\frac{T_{f}}{M_{p}}\right)^{\nu}=qT_{f}^{5}$
$\rightarrow\;\;\,T_{f}^{3-\nu}=\eta\;\frac{2\pi}{3}\sqrt{\frac{\pi g_{*}}{5}}\frac{1}{qM_{p}^{\nu+1}}.$ (36)
Following Anish , we can examine the freeze-out temperature in Barrow cosmology. The aforementioned equality allows to compute the temperature of freeze-out as
$T_{f}=M_{p}\left[\eta\;\frac{2\pi}{3}\sqrt{\frac{\pi g_{*}}{5}}\frac{1}{qM_{p}^{4}}\right]^{\frac{1}{3-\nu}}.$ (37)
Since $\delta T_{f}=T_{f}-T_{0f}$ , where $T_{0f}\sim 0.6MeV$ JohnD , by using equation (37) we have
$\left|\frac{\delta T_{f}}{T_{f}}\right|$ $=$ $\left|1-\frac{T_{0f}}{T_{f}}\right|$ (38)
$=$ $\left|1-\frac{T_{0f}}{M_{p}}\left[\eta\frac{2\pi}{3}\sqrt{\frac{\pi g_{*}}{5}}\frac{1}{qM_{p}^{4}}\right]^{-\frac{1}{3-\nu}}\right|.$
Inserting $\eta$ from Eq. (25) into the above equation, we find
$\left|\frac{\delta T_{f}}{T_{f}}\right|=\left|1-\frac{T_{0f}}{M_{p}}\left[\left(\frac{2\pi}{3}\sqrt{\frac{g_{*}}{5}}\left(\frac{2-\delta}{2+\delta}\right)^{1/{\delta}}\right)^{\frac{\delta}{2-\delta}}\right.\right.\\ \left.\left.\times\frac{2\pi}{3}\sqrt{\frac{\pi g_{*}}{5}}\frac{1}{qM_{p}^{4}}\right]^{-\frac{1}{3-\nu}}\right|,$ (39)
where
$3-\nu=\frac{6-5\delta}{2-\delta}.$ (40)
The constraint on $\delta$ parameter can be found by using Eqs. (35) and (39). In Fig. 1 we plot $\delta T_{f}/T_{f}$ from (39) vs $\delta$. The upper bound in (35) is also represented. Therefore, the constraints from BBN yield
$\delta\simeq 0.01.$ (41)
This result is consistent with constraint $0\leq\delta\leq 1$, where the parameter $\delta$ of the Barrow entropy is always bounded.
Figure 1: The behavior of $\delta T_{f}/T_{f}$ vs $\delta$ in the modified Barrow cosmology for early universe. Here $\delta T_{f}/T_{f}$ is defined in (39), while the observational upper bound of $\delta T_{f}/T_{f}$ is given by Eq. (35). Constraints from BBN requires $\delta\simeq 0.01$.
III.2 Primordial light elements ${}^{4}\textrm{He}$, D and Li in Barrow cosmology
In this section we derive the constraints on Barrow exponent $\delta$ in a different approach in which we probe the effects of Barrow cosmology on the primordial abundances of light elements, i.e., Deuterium ${}^{2}\textit{H}$, Tritium ${}^{7}\textit{Li}$, and Helium ${}^{4}\textit{He}$. We use the observational data for the abundances of aforementioned light elements to derive the bound on Barrow parameter. The main idea is to replace the amplication factor $Z(T)$ instead of the usual one entering the primordial light elements abundances, which is related to the neutrinos species effective number Luciano . We observe that in the standard cosmological model one simply has $Z=1$. Modified gravity or the presence of extra light particles such as neutrinos, may deviate Z from unity, in which case the modified Z-factor takes the form Anish ; Luciano ; Boran
$Z_{\nu}=\left[1+\frac{7}{43}(N_{\nu}-3)\right]^{1/2},$ (42)
where $N_{\nu}$ is the number of neutrinos species. The baryon-antibaryon asymmetry $\eta_{10}$, plays an important part in this analysis Adv ; Simha . Since we focus on the effects of Barrow cosmology in BBN, we set $N_{\nu}=3$, rejecting the possibility that in our framework deviations of $Z$ from unity appears due to the presence of extra particle degrees of freedom.
In what follows, we follow the approach of Anish ; Sahoo . We remind the main features:
-${}^{4}\textrm{He}$ abundance- The first step of production of ${}^{4}\textit{He}$ consists in generating ${}^{2}\textit{H}$ from a neutron and a proton. After that D is converted into Tritium and ${}^{3}\textit{He}$. The sequence of nuclear reactions that produce Helium ${}^{4}\textit{He}$ are
$n+p$ $\rightarrow\,^{2}\text{H}+\gamma,$ (43)
${}^{2}\text{H}+^{2}\text{H}$ $\rightarrow\,^{3}\text{He}+n,$ (44)
${}^{2}\text{H}+^{2}\text{H}$ $\rightarrow\,^{3}\text{H}+p.$ (45)
The Helium ${}^{4}\textit{He}$ is finally produced through the following processes
${}^{2}\text{H}+^{3}\text{H}$ $\rightarrow\,^{4}\text{He}+n,$ (46)
${}^{2}\text{H}+^{3}\text{He}$ $\rightarrow\,^{4}\text{He}+p.$ (47)
Numerical best fit constrains the primordial abundance of ${}^{4}\textit{He}$ to be Kneller ; Annu
$Y_{p}=0.2485\pm 0.0006+0.0016\left[(\eta_{10}-6)+100(Z-1)\right],$ (48)
where in our case $Z$ is given by Eq. (24). Here, the baryon density parameter $\eta_{10}$ is defined as Adv ; Simha
$\eta_{10}\equiv 10^{10}\eta_{B}\equiv 10^{10}\frac{n_{B}}{n_{\gamma}}\simeq 6,$ (49)
where $\eta_{B}\equiv{n_{B}}/{n_{\gamma}}$ is the baryon to photon ratio Wamp . For $Z=1$, we recover the standard result predicted by BBN for the ${}^{4}\textit{He}$ abundance $(Y_{p})|_{GR}=0.2485\pm 0.0006$. The observational data on ${}^{4}\textit{He}$ abundance and setting $\eta_{10}=6$ implies that $Y_{p}=0.2449\pm 0.004$ Brain . Consistency between aforementioned value of $Y_{p}$ and Eq.(48) gives
$0.2449\pm 0.0040=0.2485\pm 0.0006+0.0016\left[100(Z-1)\right].$ (50)
Thus we can fix $Z$ as
$Z=1.0475\pm 0.105.$ (51)
- ${}^{2}\textit{H}$ abundance- Deuterium is produced from the process $n+p\rightarrow\,^{2}\text{H}+\gamma$ Following the same approach as above, Deuterium abundance can be determined from the numerical best fit in Adv , giving
$y_{Dp}=2.6(1\pm 0.06)\left(\frac{6}{\eta_{10}-6(Z-1)}\right)^{1.6}.$ (52)
As before the values $Z=1$ and $\eta_{10}=6$ yield the results in standard cosmology $Y_{D_{p}}|_{GR}=2.6\pm 0.16$. Equating the observational constraint on D abundance $Y_{D_{p}}=2.55\pm 0.03$ Brain , with Eq.(52), one gets
$2.55\pm 0.03=2.6(1\pm 0.06)\left(\frac{6}{\eta_{10}-6(Z-1)}\right)^{1.6}.$ (53)
And hence the constraints on $Z$ is
$Z=1.062\pm 0.444,$ (54)
which is partially overlaps with the constraint from ${}^{4}\textit{He}$ abundance in Eq. (51).
- ${}^{7}\textit{Li}$ abundance- The $\eta_{10}$ parameter (49) on the one hand successfully fits the ${}^{4}\textit{He}$, $D$ and other light elements abundances, and on the other hand is somehow inconsistent with observed ${}^{7}\textit{Li}$ abundance. In fact, the ratio of predicted value of ${}^{7}\textit{Li}$ abundance in the standard cosmological model to the observed one lies in the range Boran
$\frac{\text{Li}|_{GR}}{\text{Li}|_{obs}}\in[2.4-4.3].$
Quite unexpectedly, neither standard BBN nor any other model are able to fit this low abundance of ${}^{7}\textit{Li}$. This puzzle is known as Lithium problem Boran . By demanding consistency between observational bounds on Lithium abundance $Y_{Li}=1.6\pm 0.3$ Brain , and the numerical best fit for ${}^{7}\textit{Li}$ abundance
$y_{Li}=4.82(1\pm 0.1)\left[\frac{\eta_{10}-3(Z-1)}{6}\right]^{2},$ (55)
where $Z$ can be fixed as
$Z=1.960025\pm 0.076675.$ (56)
This constraint does not overlap with the results for ${}^{2}\textit{H}$ abundance in Eq.(54), and ${}^{4}\textit{He}$ abundance in Eq.(51).
III.3 Discussion on the results
Let us discuss the obtained results, which we derived for the bound on the Barrow parameter $\delta$. In Fig. 2 we plot (24) taking into account the bounds presented in (51). As we can see, the parameter $\delta$ lies in the range
$0\lesssim\delta\lesssim 0.001$ (57)
Figure 2: $Z_{{}^{4}\textrm{He}}$ vs $\delta$. The experimental range is reported in (51). We have fixed $\eta_{10}=6$ and varied the temperature of freeze-out in the range $T_{f}=[0.01,10]MeV$.
We can observe the results for Deuterium (using Eq.(54)) in Fig. 3. It is seen that the corresponding range for $\delta$ is
$0\lesssim\delta\lesssim 0.008$ (58)
As we can see, the range of $\delta$ for Deuterium (58) overlaps with the range for ${}^{4}\textit{He}$ given by Eq. (57).
Figure 3: $Z_{{}^{2}\textrm{H}}$ vs $\delta$. The experimental range is reported in (54). We have fixed $\eta_{10}=6$ and varied the temperature of freeze-out in the range $T_{f}=[0.01,10]MeV$.
In the case of Lithium, Fig. 4 shows that the range of $\delta$ for ${}^{7}\textit{Li}$ does not overlap with the constraints on $\delta$ that ${}^{4}\textit{He}$ and ${}^{2}\textit{H}$ abundances provided in Eqs. (57) and (58).
Figure 4: $Z_{Li}$ vs $\delta$. The experimental range is reported in (56). We have fixed $\eta_{10}=6$ and varied the temperature of freeze-out in the range $T_{f}=[0.01,10]MeV$.
IV Relation between cosmic time and temperature
The condition of thermal equilibrium in early universe results to entropy conservation in a co-moving volume Weinberg
$s(T)a^{3}=\rm constant,$ (59)
where $s(T)$ is the entropy per unit co-moving volume. Taking the derivative of the above equation with respect to time gives
$\dot{s}(T)a^{3}+3\dot{a}a^{2}s(T)=0\mapsto\frac{ds(T)}{dt}a^{3}=-3\dot{a}a^{2}s(T).$ (60)
Using modified Friedmann equation (19), we arrive at
$\frac{ds(T)}{dt}=-3\left(\frac{8\pi G_{\rm eff}}{3}\rho\right)^{1/2-\delta}s(T).$ (61)
We can write the above equation in the form
$dt=-\frac{ds(T)}{3s(T)}\left(\frac{8\pi G_{\rm eff}}{3}\rho\right)^{1/\delta-2}.$ (62)
And hence the cosmic time $t$ can be written as
[…]
where the prime stands for the derivative with respect to temperature $T$.
In particular, during any era in which the dominant constituent of the universe is a highly relativistic ideal gas, the entropy and energy densities are given by Weinberg
$s(T)=\frac{2\mathcal{N}a_{B}T^{3}}{3}\;$ (64)
$\rho(T)=\frac{\mathcal{N}a_{B}T^{4}}{2},$ (65)
(where $\mathcal{N}$ is the number of particles and antiparticles and each spin state separately) Weinberg . Inserting $\frac{{s}^{\prime}(T)}{s(T)}=\frac{3}{T}$ and aforementioned definition for $\rho$ into the Eq. (63), we obtain
[…]
We can rewrite the above equation in the following form
[…]
Integrating yields
$t=\frac{1}{T^{4/(2-\delta)}}\left(\frac{8\pi G_{\rm eff}\mathcal{N}a_{B}}{6}\right)^{1/\delta-2}\frac{2-\delta}{4}+\rm constant.$ (68)
The above equation is the relation between the cosmic time $t$ and temperature $T$ in the context of the modified Barrow cosmology. For $\delta=0$ we find $t=\frac{1}{T^{2}}\sqrt{\frac{3}{16\pi G\mathcal{N}a_{B}}}$, which is the result of standard cosmology Weinberg .
Let us note that we have used the modified Friedmann equation which has been derived in the units where $\hbar=k_{B}=c=1$. On the other hand, the right hand side of Eq. (68) should have the dimension $[\,t\,]$. If we take into account all the constants in deriving the modified Friedmann equations, the effective gravitational constant in Eq. (18) takes the form
$G_{\mathrm{eff}}\equiv\frac{A_{0}}{4}\left(\frac{2-\delta}{2+\delta}\right)\left(\frac{A_{0}}{4\pi}\right)^{\delta/2}{c^{2-\delta}\over\alpha_{B}\Gamma_{B}},$ (69)
where we have defined
$\alpha_{B}\equiv k_{B}\left(c^{3}\over\hbar\right)^{1+(\delta/2)}\quad\,,\,\,\quad\Gamma_{B}\equiv{\hbar c\over k_{B}}.$ (70)
Thus, one can easily check that the right side of Eq. (68) has dimension $[\,t\,]$. In early universe, when the constituents of the universe were photons, plus three species of neutrinos and three of antineutrinos, plus electrons and positrons, we have $\mathcal{N}=43/4$, so that Eq. (68) yields, in cgs units,
$t=\left({T\over 10^{10}K}\right)^{4\over\delta-2}F(\delta)+\rm{constant}.$ (71)
Where we have defined
$F(\delta)$ $=$ $\left[\left(\frac{2-\delta}{2+\delta}\right)\times(9.2\times 10^{-88})^{\delta/2}\right]^{1\over\delta-2}$ (72)
×$\times{2-\delta\over 4}(0.278)^{1/(\delta-2)}$
Again for $\delta=0$ we have $F(\delta)\to 0.994\;sec$ and the result of standard cosmology is restored Weinberg
[…]
In Fig. 5 we plot the behavior of the temperature $T$ as a function of cosmic time $t$ for different values of Barrow parameter $\delta$. We observe that as the fractal structure increases, the temperature of the early universe increases too.
Figure 5: The behavior of $T$ vs $t$ in modified Barrow cosmology for early universe.
V Closing remarks
In this paper we have used BBN data in order to find the bounds on the free parameter $\delta$ of Barrow entropy, which is an extended entropy relation originating from the incorporation of quantum gravitational effects on the structure of a black hole, parameterized by $\delta$. Barrow entropy leads to modified Firedmann equations containing extra terms. These extra terms should be sufficiently small in order not to spoil BBN epoch. We can summarize our results in this paper as follows:
(i) We first presented an analysis of the BBN epoch in the Barrow cosmology and calculated the deviations of the temperature of freeze-out in comparison to the result of standard cosmology. We have used observational upper bound on $\left|\frac{\delta T_{f}}{T_{f}}\right|$ which is the allowed deviation from GR, in order to extract the bound on Barrow exponent $\delta$. We have shown that the BBN constraint imposes a value for the Barrow exponent $\delta$ of the order $\delta\simeq\mathcal{O}(0.01)$. As expected this result shows that the deformation from Bekenstin-Hawking expression should be small.
(ii) We have also explored the consequences of Barrow cosmology on the primordial light elements formation. The analysis of the light elements ${}^{4}\textit{He}$, ${}^{2}\textit{H}$,${}^{7}\textit{Li}$ shows that the range of variability of $\delta$, using the bounds (51), (54) and (56), does not overlap.
(iii) We have presented the relation between cosmic time $t$ and the temperature $T$ of the universe in the modified Barrow cosmology. We have then plotted the behavior of temperature as a function of cosmic time $t$ for different $\delta$. The result shows that as the fractal structure increases, the temperature of the early universe increases, too.
Following the present study, we can use BBN data in order to impose constraints on the free parameter of other entropy corrected cosmological models such as modified cosmology inspired by zero-point length, or the modified Kaniadakis cosmology. We can also investigate the relation between time and temperature in these models. We leave these for the future investigations.
Acknowledgements.
We thank Shiraz university Research Council. The work of A. Sheykhi is based upon research funded by Iran National Science Foundation (INSF) under project No. 4022705.